AnimeGhoul8863:

The coordinate plane below represents a town. Points A through F are farms in the town. graph of coordinate plane. Point A is at 2, negative 3. Point B is at negative 3, negative 4. Point C is at negative 4, 2. Point D is at 2, 4. Point E is at 3, 1. Point F is at negative 2, 3. Part A: Using the graph above, create a system of inequalities that only contains points C and F in the overlapping shaded regions. Explain how the lines will be graphed and shaded on the coordinate grid above. (5 points) Part B: Explain how to verify that the points C and F are solutions to the system of inequalities created in Part A. (3 points) Part C: Chickens can only be raised in the area defined by y < 5x − 3. Explain how you can identify farms in which chickens can be raised. (2 points)

AnimeGhoul8863:

1 attachment
Ultrilliam:

@dude Save this person

AnimeGhoul8863:

he is eating dinner i just want to finish

Ultrilliam:

ah @563blackghost

AnimeGhoul8863:

do you think u can help

Bearclaws72:

Out of my skill level sorry bud

AnimeGhoul8863:

OMG NO ONE CAN HELPME

AnimeGhoul8863:

this site is surpost to be full of people who want to help

AnimeGhoul8863:

ill ask more people

AnimeGhoul8863:

PLEASE HELP ME

AnimeGhoul8863:

no one knows 9th grade math

AnimeGhoul8863:

Finally can u help

563blackghost:

So sorry I left QC open (cause of the mining) I had left to go get food ;-;

AnimeGhoul8863:

its ok just please tell me u can help or im doomed

563blackghost:

i can :) just gotta find an inequality to fit

AnimeGhoul8863:

i thought of f>g

AnimeGhoul8863:

and do you think we will get it done Like in the next 10 mins

AnimeGhoul8863:

ultrilliam can she break the rules just this once and give me a direct answer im only asking u bc ur the founder of the website

563blackghost:

sorry i can't do that...

AnimeGhoul8863:

ok but can we try to get it done quick its late and im tired and want to finish

AnimeGhoul8863:

Hello we can start on part a and go from there

Angle:

we want two inequalities that include the points c and f

Angle:



AnimeGhoul8863:

i said F>G

Angle:

what are the equations of the red lines I drew, that would be the answer to part A

AnimeGhoul8863:

(4,-4) and (5,-5)

Angle:

I

AnimeGhoul8863:

ummmmmmm idk f(x)4+?

AnimeGhoul8863:

Hello

Angle:

you know y = mx + b if the line passes through the points (0, 4) and (-4, 0) what is the equation of the line?

AnimeGhoul8863:

y=4x+-4

Angle:

Not quite! y = mx + b where m is the slope and b is the y intercept now, what is the slope between the points (0, 4) and (-4, 0) ?

AnimeGhoul8863:

-4

Angle:

the equation for slope is: rise / run or \(\huge \frac{y_2 - y_1}{x_2 - x_1}\)

AnimeGhoul8863:

ok and

AnimeGhoul8863:

OMG PLEASE GIVE ME A DIRECT ANSWER IM DIEING HERE I JUST WANT TO FINISH

AnimeGhoul8863:

11 hours of test is too much for a person

Angle:

\(\huge \frac{0-4}{-4 - 0} = ~?\)

AnimeGhoul8863:

1

AnimeGhoul8863:

am i right

dude:

Yes

AnimeGhoul8863:

ok now what

Angle:

So what is the equation for the line that passes through the points (0, 4) and (-4, 0)?

AnimeGhoul8863:

idk 1

dude:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Angle Not quite! y = mx + b where m is the slope and b is the y intercept now, what is the slope between the points (0, 4) and (-4, 0) ? \(\color{#0cbb34}{\text{End of Quote}}\)

dude:

What is the b value?

AnimeGhoul8863:

y=-4+1

dude:

https://www.desmos.com/calculator/cryojfbdzu This is what we have, hot do you have the line to pass through both points? Feel free to change the equation there

AnimeGhoul8863:

i still DONT GET IT!!!!!!!

dude:

What is the y-intercept?

AnimeGhoul8863:

0,4

Angle:

perfect then what is b equal to? b = the y intercept, when x = 0 for example if the y intercept is (0, 3) then b = 3 so for this problem, if the y intercept is (0, 4) b = ?

AnimeGhoul8863:

i said that already y=-4+4 and u made me do all this

AnimeGhoul8863:

I JUST NEED TO FINISH ITS 8:30 AT NIGHT

dude:

We stated that the slope was 1 and the y-inetercept was 4, just substitute those values into \[y=mx+b\\ \color{red}{\large{y=1x+4}}\]

AnimeGhoul8863:

y=-4x+4 is the equation

AnimeGhoul8863:

so its 1x+4

dude:

-gbSpT-MTUSnL1qG1rQKhQ.png Yes

AnimeGhoul8863:

so y=1x+4 is part a whats part b

Angle:

We're not done with part A They want a SYSTEM of equations which means more than one equation

AnimeGhoul8863:

IDFK

Angle:

First, we want to shade in the area above the equation we just found would we use < , > , or = to describe the area we shade above the line?

AnimeGhoul8863:

<

AnimeGhoul8863:

ok i got 10 mins till my time is up we need to hurry im begging you

Angle:

Why do you think it is < "less than" ? we want the equation that is shaded above the line right? so what is the symbol for "greater than" ?

AnimeGhoul8863:

i dont know i have 10 FREAKING MINS HELP ME QUICK STOP ASKING 3 QUESTION u wont give a direct answer fine but be fast with the answering PLEASE

AnimeGhoul8863:

i cant take the waiting any more

Ultrilliam:

There's a old saying, don't bite the hand that feeds you :|

AnimeGhoul8863:

oh i will bc i will fail if i dont

Angle:

the "greater than" symbol is > Because of this fact, the equation becomes y > x + 4

Angle:

the next equation passes through the points (0, 7) and (-7, 0) using what we have learned a few minutes ago, find the equation of that line

Angle:

oh did you just say that the equation is y = x + 7 WOW how very smart of you I'm so glad you were able to do that part yourself then next we want to know about the part shaded UNDER that line that would be less than right? what is the symbol for "less than"? did you say < fantastic! so our equation would be y < x + 7 now can you repeat to me what are the two equations we have found so far?

AnimeGhoul8863:

y<1+4

AnimeGhoul8863:

i mean y=1+4 and y<x+7

Angle:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Angle the "greater than" symbol is > Because of this fact, the equation becomes y > x + 4 \(\color{#0cbb34}{\text{End of Quote}}\)

AnimeGhoul8863:

so its > not <

Angle:

could you repeat for me what you think the answer for Part A is? I want to make sure you have the correct answer

AnimeGhoul8863:

y=1+4 and y<x+7

Angle:

hmm y = 1 + 4 is not quite right \(\color{#0cbb34}{\text{Originally Posted by}}\) @Angle the "greater than" symbol is > Because of this fact, the equation becomes y > x + 4 \(\color{#0cbb34}{\text{End of Quote}}\)

AnimeGhoul8863:

Chicken GIVE THE F-ing answer or im leaving i cant take the beat around the bush anymore

Angle:

did you say that you think the answer to part A is y > x + 4 and y < x + 7 ? WOW you are so smart I'm so proud of you for getting that answer without my help now we can finally move onto part b

AnimeGhoul8863:

thank you

Angle:

to prove that some point is a part of the system of equations we just found, we should plug that point into the equations and see if it comes out to be true! so the point for C is (-4, 2) what happens when we plug that into y > x + 4 and y < x + 7 ? . . . did you say that y > x + 4 becomes 2 > -4 + 4 which is simplified to 2 > 0 awesome just look at how much you are learning from this magnificent experience now we can see that 2 > 0 is a TRUE statement then y < x + 7 becomes 2 < -4 + 7 which is simplified to 2 < 3 holy cows this is also a true statement! this means that these the point C works for these two equations that we found! WOW AMAZING

Angle:

now that we proved it for point C, we also have to prove it for point F as well! can you do that? what happens when you plug in (-2, 3) for (x, y) into our equations??? ... y > x + 4 becomes 3 > -2 + 4 which is simplified to 3 > 2 which is a TRUE STATEMENT WHOA that's amazing! let's see if it is right for the other equation too! y < x + 7 becomes 3 < -2 + 7 which is simplified to 3 < 5 this is also a TRUE STATEMENT and you did all of this math yourself? without my help wow look at how much smarter you've gotten no complaining or anything so that's how you proved that both points C and F are in the equations we found? awesome

dude:

@AnimeGhoul8863 you alive?

AnimeGhoul8863:

yes i was reading it

AnimeGhoul8863:

thx

AnimeGhoul8863:

gtg

Ultrilliam:

Hope to see you again soon :|

Angle:

I won't be helping you with part C then have a nice day good luck with the rest of your homework and stuff it was nice teaching you so much I'm glad you learned a lot with me